Circular No. 6840

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

AX J0049–729 = RX J0049.1–7250 P. Kahabka, University of Amsterdam, and W. Pietsch, Max-Planck-

Institut für Extraterrestrische Physik, communicate: “The new x-ray pulsar in the Small Magellanic Cloud reported by Yokogawa and Koyama (IAUC 6835) coincides in position (within the 2′ error circle) with the ROSAT source RX J0049.1–7250 (Kahabka and Pietsch 1996, A.Ap. 312, 919). The source is highly variable and has been detected during ROSAT PSPC observations on 1991 Oct. 9 and 1992 Apr. 15–24, but it was not detected during a 35 000-s ROSAT HRI observation on 1997 May 9–25. A mean source position derived from reprocessed archival ROSAT data is α = 0^{h}49^{m}02^{s}.3, δ = −72^{o}50′55′′ (equinox 2000.0; ± 13′′ at 90-percent confidence). The source may be a Be-type transient. Optical identification work is highly encouraged.”

SAX J1818.6–1703 AND KS 1741–293 J. in ’t Zand and J. Heise, Space Research Organization of the Nether-

lands (SRON), Utrecht; M. Smith, BeppoSAX Science Operation Center, Rome, and SRON; J. M. Muller, BeppoSAX Science Data Center, Rome, and SRON; and P. Ubertini and A. Bazzano, Istituto di Astrofisica Spaziale, CNR, Rome, report: “In an observation of the galactic bulge with Wide Field Camera 2 onboard BeppoSAX during Mar. 10.2–12.0 UT, two interesting x-ray sources were detected other than SAX J1810.8–2609 (IAUC 6838). One was seen to turn on at Mar. 11.8 and peak at Mar. 11.86 at a level of 0.1 Crab in 2–9 keV and 0.4 Crab in 9–25 keV (as measured over 15-min time bins). Over the last 40 min of the observation, the intensity is at least twice as low. The 3σ upper limits for the pre-flare part of the 2-day observation are 10 and 35 mCrab in these respective energy bands. The position is α = 18^{h}18^{m}39^{s}, δ = −17^{o}03′.1 (equinox 2000.0, preliminary error radius 3′). No catalogued x-ray source is within the error box, and we designate this new transient SAX J1818.6–1703. The error box contains one catalogued optical star, HD 168078 (a B3 III star with m_{v} = 10.7). Another hard source was detected at α = 17^{h}44^{m}40^{s}, δ = −29^{o}20′.5 (preliminary error radius 3′). There is one catalogued x-ray source in the error box, KS 1741–293 (a burst source first identified in 1989 and not detected since). Intensities are 6 mCrab in 2–9 keV and 25 mCrab in 9–25 keV. The intensity appears constant within ∼ 30 percent on a timescale of half a day. Further observations are encouraged for both sources.”

1998 March 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green